# D-69120 Heidelberg, Germany

### v2

@@ Abstract
Aims. Understanding the rotational dynamics of interstellar dust grains is quintessential for the analysis of the observed dust polarization signal. Additionally, high rotation velocities may cause rotational disruption of the grains, impacting the grain size distribution. We aim to constrain the set of parameters for an accurate description of the rotational spin-up process of ballistic dust grain aggregates driven by radiative torques (RATs).
Methods. We model the dust grains as complex fractal aggregates grown by the ballistic aggregation of uniform spherical particles (monomers) of different sizes. A broad variation of dust materials, shapes, and sizes are studied in the presence of different radiation sources.
Results. We find that the canonical parametrization for the torque efficiency overestimates the maximal angular velocity $\omega_{\mathrm{RAT}}$ caused by RATs acting on ballistic grain aggregates, and to resolve this problem we propose a new parametrization which predicts $\omega_{\mathrm{RAT}}$ more accurately. We find that RATs are most efficient for larger grains with smaller monomer-density, this manifests as size and monomer-density dependence in the constant part of the parametrization. Following the constant part, the parametrization has two power-laws with different slopes, which retain the universality for all grain sizes. The maximum grain rotation does not scale linearly with radiation strength, due to different drag mechanisms dominating depending on grain material and environment. The angular velocity $\omega_{\mathrm{RAT}}$ of individual single dust grains has a wide distribution and may even differ from the mean by up to two orders of magnitude. Even though ballistic aggregates have a lower RAT efficiency, strong sources of radiation (larger than $\approx 100$ times the typical interstellar radiation field) may still produce rotation velocities large enough to cause rotational disruption of dust grains.
@@ 1 Introduction
Dust is a common component in astronomic environments, from galaxies to stellar discs and therefore plays a crucial role in observations. Most astronomers come into contact with dust because of dust extinction, which affects almost all observations (Draine 2003). Calzetti et al. (2000) as well as Bernstein et al. (2002) estimated, that $30\ \%$ or more of the energy produced by stars in the universe is absorbed by dust and re-emitted in the infrared. The dust helps to cool down gas, facilitating gravitational collapse (Omukai et al. 2005), also playing a major role in later stages of star formation, as well as constituting the first stage in the formation of planets (Dullemond et al. 2008; Liu & Ji 2020). Rapidly rotating dust grains are the common source for polarized light, carrying information about the magnetic field orientation (see e.g. Lazarian 2007; Andersson et al. 2015, for review). The catalysis of molecular hydrogen (Hollenbach & Salpeter 1971; Cazaux & Tielens 2002) and other astrochemical reactions in the ISM (Dishoeck 2014) is also impacted by grain rotation (Hoang & Tram 2020). The first models trying to explain the observed dust alignment, like paramagnetic relaxation (Davis & Greenstein 1951) or the Gold mechanism (Gold 1952) both encountered problems and were ultimately unable to account for the observed alignment. A big leap in the study of alignment was made, when Purcell (1979) realized, that grains could experience systematic torques acting in a specific direction in grain coordinates for large time spans compared to the timescale it takes gas collisions to randomize the grain rotation. With such systematic torques the grain can spin up to much faster angular velocities than the typical thermal rotation rate. With suprathermal rotation, a grain can remain in a given orientation for much longer than the gas collision timescale, allowing alignment of angular momentum with the maximal moment of inertia via inelastic relaxation or the Barnett relaxation discovered by Purcell (1979), based on the Barnett (1917) effect, which causes magnetization along the rotation axis in a paramagnetic body. If the grains rotation axis is not aligned with a principal axis of inertia, the rotation axis and therefore the magnetization axis will precess in body coordinates, dissipating energy and aligning the rotation axis with the axis of maximal moment of inertia. In the presence of an external magnetic field the Davis-Greenstein-process can align the grain with the magnetic field (Purcell 1979).
@@
A one to one mapping of dust alignment to magnetic field lines would of course be of great use to the study of magnetic fields. Polarization observations of the object in question could then always be used to obtain information about the magnetic field morphology. However uncertainties about the completeness of the model remained e.g. no increase in polarization with extinction in cold dark clouds (Goodman et al. 1995) and alignment of grains in cometary dust (Harwit & Vanýsek 1971), so the exploration of other alignment mechanisms continued.
@@
The idea of systematic radiative torques (RATs) as the cause for rapid grain rotation had already been advanced decades prior (Dolginov 1972; Dolginov & Mytrophanov 1976), but only became an integral part of dust alignment theory with the work of Draine & Weingartner (1996). Using the discrete dipole approximation (DDA) method it was shown, that scattering and absorption of light from irregular grains results in torques acting on timescales much larger than random gas collisions. Their calculations were done on a irregular shape of intersecting cuboids and it was found, that radiative torques depend strongly on grain size. Additional grain shapes were investigated in the following work (Draine & Weingartner 1997) in the context of anisotropic radiation, $\mathrm{H}_{2}$-formation torques and paramagnetic relaxation. In most cases the grain reaches long-term stable conditions by aligning the grains principal axis with the magnetic field. Later, Lazarian & Hoang (2007a) (LH07 hereafter) introduced an analytic model (AMO) for a spherical grain attached to a flat rectangular mirror via a weightless non interacting rod in this way capturing the helical properties of irregular grains. The AMO was compared to DDA simulations of irregular grain models, corresponding to the grain shapes in Draine & Weingartner (1997) and found a good agreement. The alignment of grains principal axis with the direction of incoming radiation was confirmed when magnetic field effects were subdominant. The parametrization of the RAT spin-up efficiency was found to be a power-law with an exponent of about $-3$. Physical implications of RATs were explored in following studies (Lazarian & Hoang 2008; Hoang & Lazarian 2008, 2009, 2016) and the model was expanded by the addition of superparamagnetic relaxation due to iron inclusions. This increased the range of parameters in which grains can efficiently align with the magnetic field which is necessary to explain high degrees of polarization from observations (Planck Collaboration et al. 2020). The AMO was used successfully to model synthetic dust polarization observations, as long as free alignment parameters could be used to adjust the model (Bethell et al. 2007; Hoang & Lazarian 2014; Reissl et al. 2016). Herranen et al. (2019) then made a step towards testing more realistic grain shapes: gaussian ellipsoids, oblate spheroids, prolate spheroids, and spheres with random deformations and found a shallower slope of $-2.6$ for the power-law for the RAT parametrization. In this work we take one more step towards more realistic grain shapes, by simulating random typically fractal aggregate grains composed of different sized monomers and compare their RATs to previous results. The motivation is to model the grain growth ab-initio using the current understanding, that dust is originally formed in AGB stellar winds or supernovae (Todini & Ferrara 2001; Schneider et al. 2014; Nanni et al. 2013). As the dust forming materials in the winds cool down, heavy elements condense into spherical particles - in the following called monomers - which are the building blocks for larger dust grains formed by monomer accretion and coagulation of grains (Gobrecht et al. 2016; Norris et al. 2012). More realistic grain modeling may result in differing interactions with radiation. One example being the study of Arnold et al. (2019), that found significant differences in dust blowout sizes between compact and agglomerated grains.
@@
The aim of this work is to improve the understanding of RATs by simulating individual grains composed of monomers for different sources of radiation. The properties of grains and the way they are modeled is presented in Sect. 2. The theoretical basis for our simulations are presented in Sect. 3 and models for the sources in Sect. 4, usage of the simulation codes is explained in Sect. 5, results are shown in Sect. 6 and discussed in Sect. 7, a summary and outlook are given in Sect. 8.
@@ 2 Ballistic dust grain aggregation
In this work dust grains are modeled as conglomerates made up of spherical particles of a single material, called monomers. The monomers stay attached to each other due to the intermolecular forces at their contact area (see e.g. Johnson et al. 1971). The growth of larger dust grains follows an aggregation process by means of ballistic collision and sticking of individual monomers (Dominik & Tielens 1997). To mimic the grain growth process of ballistic aggregation (BA) we follow the procedure as outlined in Shen et al. (2008) modified to work with monomers of variable sizes (Reissl et al. 2023). Starting with two or three connected monomers new monomers with radii drawn from a log-normal distribution between $10\ \mathrm{nm}$ and $100\ \mathrm{nm}$ are injected into the simulation volume from random directions, sticking to the grain upon collision.
@@
Grains with an effective radius $a_{\mathrm{eff}}$ [^1] The effective grain radius gives the radius of the grain if it was reshaped into a perfect sphere. from $50\,\mathrm{nm}$ up to $250\,\mathrm{nm}$ in steps of $50\,\mathrm{nm}$ are created. To reach an exact effective radius the sizes of the final monomers added to the grain are sampled in a biased way (lowering the upper bound on monomer size), until the desired $a_{\mathrm{eff}}$ is exactly reached. In addition to BA where each monomer remains at its initial sticking position, we model different monomer-densities or porosities of aggregates by means of monomer migration. After hitting the forming grain the added monomer rolls along the surface of the monomer it impacted until it touches a second monomer of the grain (BAM1). Here, all monomers are connected to at least two other monomers. For the BAM2 case each monomer rolls twice until it touches a third monomer. For BAM2, all monomers are connected to at least three other monomers. This is the most dense a random grain can be packed with monomers using the migration process. It follows, that BA-grains are the least dense/most porous and BAM2-grains are the most dense/least porous, an exemplary grain of each monomer-density is shown in Fig. 1.
@@
To recapitulate the three grain properties are: the material from which the grain is made (astrosilicate or hydrogenated amorphous carbon) which differ in their optical properties, the monomer-density of the grain (BA,BAM1,BAM2) determining how closely monomers are packed and finally grain size given by the effective radius ($50\,\mathrm{nm}$ to $250\,\mathrm{nm}$ in steps of $50\,\mathrm{nm}$). For every combination of these three properties 30 different grains are created, leading to a total of 900 individual grains. This method of modeling ballistic grain aggregates was previously used in Reissl et al. (2023).
@@
Figure 1: Exemplary BA (left), BAM1 (middle), BAM2 (right) grains, respectively, with an effective radius of ${a_{\mathrm{eff}}=250\ \mathrm{nm}}$. The rotation axis $\hat{a}_{\mathrm{1}}$ is depicted as a red arrow and is associated with the direction of the maximal moment of inertia $I_{\mathrm{max}}$.
3 Maximal grain rotation
@@
In the presence of a directed radiation field, irregularly shaped dust grains can experience momentum transfer from photon-interactions. In general the time averaged forces acting on such a grain may result in a net-torque, accelerating the rotation. This radiative torque (RAT) can be evaluated by (see e.g. Draine & Weingartner 1996, and LH07)
@@
where $\gamma$ describes the anisotropy of the radiation field, $\lambda$ is the wavelength and the quantity $u_{\lambda}=F_{\lambda}/c$ is the spectral energy density of a radiation field with the spectral flux $F_{\lambda}$, whereas $c$ is the speed of light. These factors determine the overall strength of the torque independent of grain shape and orientation.
The effect of the individual grain shape and material is contained within the dimensionless spin-up efficiency vector ${Q}_{\Gamma}$ of the RAT. A perfectly spherical grain would experience no net torque at all, due to the symmetry between individual collisions resulting in clockwise or anticlockwise torques. However, for grains with irregular shapes, the contributions of individual radiation-dust interactions in a nonzero ${Q}_{\Gamma}$ vector, which modulates the orientation i.e. the sign as well as the magnitude of ${\Gamma}_{\mathrm{RAT}}$.
In previous works a power-law parametrization for $Q_{\Gamma}=|\@vec{Q}_{\Gamma}|$ was suggested (LH07)
@@
In this paper we refer to this relation as the canonical parametrization. Here, the constant part transitions into a power-law at wavelengths equal to $1.8$ times the effective radius of the grain. The power-law exponent typically lies between $\alpha=-2.6$ (Herranen et al. 2019) and $\alpha=-4$ (LH07).
For our grain dynamics modeling we assume stable alignment of the grains rotation with the direction of incoming radiation. For a single grain stably rotating about its principal axis the maximal possible angular momentum $\omega_{\mathrm{RAT}}$ caused by RATs can be calculated as
@@
where $\tau_{\mathrm{drag}}$ is the drag-timescale and $I_{\mathrm{max}}$ is the largest moment of inertia, defining the $a_{1}$-axis in the body-centric system of the grain (see e.g. Draine & Weingartner 1996, as well as Fig.  1). The total drag acting on the grain is composed of the collisions with the surrounding gas and the loss of angular momentum by emitting infrared photons. The total drag-timescale can be evaluated from the gas-drag-timescale $\tau_{\mathrm{gas}}$ and photon-drag-timescale $\tau_{\mathrm{IR}}$ with
$\tau_{\mathrm{drag}}=\frac{1}{\tau_{\mathrm{gas}}^{-1}+\tau_{\mathrm{IR}}^{-1}}$ (4)
where a shorter timescale results in a stronger drag. Here, the gas-drag-timescale
@@
The photon-drag-timescale
$\tau_{\mathrm{IR}}=\frac{c^{2}I_{\mathrm{max}}}{a_{\mathrm{eff}}^{2}f_{\mathrm{IR}}(T_{\mathrm{dust}})}\quad,$ (6)
depends on the integral over the photon emission
@@
Here, $B_{\lambda}\left(T_{\mathrm{dust}}\right)$ is the Planck function modulated by the dimensionless efficiency of absorption $Q_{\mathrm{abs}}(\lambda)$, where as $T_{\mathrm{dust}}$ is the dust temperature. For simplicity, the dust is illuminated by a single source without extinction. For the dust heating we assume thermal equilibrium. Therefore, the temperature of a single dust grain can be evaluated via
$\int_{0}^{\infty}F_{\lambda}Q_{\mathrm{abs}}(\lambda)\,\mathrm{d}\lambda=4\pi\int_{0}^{\infty}B_{\lambda}\left(T_{\mathrm{dust}}\right)Q_{\mathrm{abs}}(\lambda)\,\mathrm{d}\lambda\,,$ (8)
by equating the source flux $F_{\lambda}$ absorbed by the grain with the characteristic emitted flux $B_{\lambda}Q_{\mathrm{abs}}(\lambda)$ of each grain.
@@
 [^2] Includes: stellar radius $R$ and effective temperature $T_{\mathrm{eff}}$ for the stellar sources S1 and S2, bolometric luminosity $L_{\mathrm{bol}}$ for all localized sources, anisotropy $\gamma$ and radiation energy density $u$ given in relation to $u_{\mathrm{ISRF}}=8.64\cdot 10^{-14}\ \mathrm{J\ m^{-3}}$. The distance to the localized sources S1, S2 and AGN, is taken to be $10\,\mathrm{pc}$.
Figure 2: Spectral flux of sources ISRF, S1, S2, and AGN, respectively, utilized to irradiate the grain aggregates.

### v3

@@ Abstract
Aims. It is quintessential for the analysis of the observed dust polarization signal to understand the rotational dynamics of interstellar dust grains. Additionally, high rotation velocities may rotationally disrupt the grains, which impacts the grain-size distribution. We aim to constrain the set of parameters for an accurate description of the rotational spin-up process of ballistic dust grain aggregates driven by radiative torques (RATs).
Methods. We modeled the dust grains as complex fractal aggregates grown by the ballistic aggregation of uniform spherical particles (monomers) of different sizes. A broad variation of dust materials, shapes, and sizes were studied in the presence of different radiation sources.
Results. We find that the canonical parameterization for the torque efficiency overestimates the maximum angular velocity $\omega_{\mathrm{RAT}}$ caused by RATs acting on ballistic grain aggregates. To resolve this problem, we propose a new parameterization that predicts $\omega_{\mathrm{RAT}}$ more accurately. We find that RATs are most efficient for larger grains with a lower monomer density. This manifests itself as a size- and monomer-density dependence in the constant part of the parameterization. Following the constant part, the parameterization has two power laws with different slopes that retain universality for all grain sizes. The maximum grain rotation does not scale linearly with radiation strength because different drag mechanisms dominate, depending on the grain material and environment. The angular velocity $\omega_{\mathrm{RAT}}$ of individual single dust grains has a wide distribution and may even differ from the mean by up to two orders of magnitude. Even though ballistic aggregates have a lower RAT efficiency, strong sources of radiation (stronger than $\approx 100$ times the typical interstellar radiation field) may still produce rotation velocities high enough to cause the rotational disruption of dust grains.
@@ 1 Introduction
Dust is a common component in astronomic environments from galaxies to stellar disks. It therefore plays a crucial role in observations. Most astronomers come into contact with dust because of dust extinction, which affects almost all observations (Draine 2003). Calzetti et al. (2000) and Bernstein et al. (2002) estimated that $30\ \%$ or more of the energy that is produced by stars in the Universe is absorbed by dust and is reemitted in the infrared. The dust helps to cool the gas, which facilitates gravitational collapse (Omukai et al. 2005) and also plays a major role in later stages of star formation. It also constitutes the first stage in the formation of planets (Dullemond et al. 2008; Liu & Ji 2020). Rapidly rotating dust grains are the common source of polarized light and carry information about the magnetic field orientation (see e.g. Lazarian 2007; Andersson et al. 2015, for review). The catalysis of molecular hydrogen (Hollenbach & Salpeter 1971; Cazaux & Tielens 2002) and other astrochemical reactions in the interstellar medium (ISM) (Dishoeck 2014) is also impacted by grain rotation (Hoang & Tram 2020). The first models trying to explain the observed dust alignment, that is, paramagnetic relaxation (Davis & Greenstein 1951) or the Gold mechanism (Gold 1952), both encountered problems and were ultimately unable to account for the observed alignment. A great leap in the study of alignment was made when Purcell (1979) realized that grains could experience systematic torques acting in a specific direction in the grain coordinates for long time spans compared to the timescale that it takes gas collisions to randomize the grain rotation. With these systematic torques, the grain can spin up to much faster angular velocities than the typical thermal rotation rate. With suprathermal rotation, a grain can remain in a given orientation for much longer than the gas-collision timescale. This allows an alignment of the angular momentum with the maximum moment of inertia via inelastic relaxation or the Barnett relaxation, which was discovered by Purcell (1979) and is based on the Barnett (1917) effect, which causes magnetization along the rotation axis in a paramagnetic body. When the grain rotation axis is not aligned with a principal axis of inertia, the rotation axis and therefore the magnetization axis will precess in body coordinates, dissipating energy and aligning the rotation axis with the axis of the maximum moment of inertia. In the presence of an external magnetic field, the Davis-Greenstein process can align the grain with the magnetic field (Purcell 1979).
@@
A one-to-one mapping of dust alignment to the magnetic field lines would be of great use to the study of magnetic fields. Polarization observations of the object in question could then always be used to obtain information about the magnetic field morphology. Uncertainties about the completeness of the model remained, however, such as the lack of an increase in polarization with extinction in cold dark clouds (Goodman et al. 1995) and the alignment of grains in cometary dust (Harwit & Vanýsek 1971). The exploration of other alignment mechanisms therefore continued.
@@
The idea of systematic radiative torques (RATs) as the cause for rapid grain rotation had already been advanced decades before (Dolginov 1972; Dolginov & Mytrophanov 1976), but it only became an integral part of dust alignment theory with the work of Draine & Weingartner (1996). Using the discrete dipole approximation (DDA) method, the authors showed that scattering and absorption of light from irregular grains results in torques that act on timescales that are much longer than random gas collisions. The calculations were made on a irregular shape of intersecting cuboids, and the authors found that radiative torques strongly depend on the grain size. Additional grain shapes were investigated in the following work (Draine & Weingartner 1997) in the context of anisotropic radiation, $\mathrm{H}_{2}$-formation torques, and paramagnetic relaxation. In most cases, the grain reached long-term stable conditions by aligning the principal grain axis with the magnetic field. Lazarian & Hoang (2007a) (LH07 hereafter) later introduced an analytic model (AMO) for a spherical grain that is attached to a flat rectangular mirror via a weightless noninteracting rod. In this way, they captured the helical properties of irregular grains. The AMO was compared to DDA simulations of irregular grain models that corresponded to the grain shapes in Draine & Weingartner (1997), and a good agreement was found. The alignment of the principal grain axis with the direction of incoming radiation was confirmed when magnetic field effects were subdominant. The parameterization of the RAT spin-up efficiency was found to be a power law with an exponent of about $-3$. Physical implications of RATs were explored in following studies (Lazarian & Hoang 2008; Hoang & Lazarian 2008, 2009, 2016), and the model was expanded by the addition of superparamagnetic relaxation due to iron inclusions. This increased the range of parameters in which grains can efficiently align with the magnetic field, which is necessary to explain high degrees of polarization from observations (Planck Collaboration et al. 2020). The AMO was used successfully to model synthetic dust polarization observations as long as free alignment parameters could be used to adjust the model (Bethell et al. 2007; Hoang & Lazarian 2014; Reissl et al. 2016). Herranen et al. (2019) then made a step toward testing more realistic grain shapes: They used Gaussian ellipsoids, oblate spheroids, prolate spheroids, and spheres with random deformations and found a shallower slope of $-2.6$ for the power law for the RAT parameterization. In this work, we take one more step toward more realistic grain shapes by simulating random typically fractal aggregate grains composed of differently sized monomers. We compare their RATs to previous results. The motivation is to model the grain growth ab initio using the current understanding that dust is originally formed in AGB stellar winds or supernovae (Todini & Ferrara 2001; Schneider et al. 2014; Nanni et al. 2013). As the dust-forming materials in the winds cool down, heavy elements condense into spherical particles (in the following called monomers), which are the building blocks for the larger dust grains that are formed by monomer accretion and coagulation of grains (Gobrecht et al. 2016; Norris et al. 2012). A more realistic grain modeling may result in differing interactions with radiation. One example was given by Arnold et al. (2019), who found significant differences in dust blowout sizes between compact and agglomerated grains.
@@
The aim of this work is to improve the understanding of RATs by simulating individual grains composed of monomers for different sources of radiation. The properties of grains and the way they are modeled is presented in Sect. 2. The theoretical basis for our simulations is presented in Sect. 3, and models for the sources are described in Sect. 4. The simulation codes are explained in Sect. 5, the results are shown in Sect. 6 and discussed in Sect. 7, and a summary and outlook are given in Sect. 8.
@@ 2 Ballistic dust grain aggregation
We modeled dust grains as conglomerates consisting of spherical particles of a single material, called monomers. The monomers stay attached to each other due to the intermolecular forces at their contact area (see e.g. Johnson et al. 1971). The growth of larger dust grains follows an aggregation process by means of ballistic collision and sticking of individual monomers (Dominik & Tielens 1997). To mimic the grain-growth process of ballistic aggregation (BA), we followed the procedure as outlined in Shen et al. (2008), modified to work with monomers of variable sizes (Reissl et al. 2023). Starting with two or three connected monomers, new monomers with radii drawn from a log-normal distribution between $10\ \mathrm{nm}$ and $100\ \mathrm{nm}$ were injected into the simulation volume from random directions so that they stuck to the grain upon collision.
@@
Grains with an effective radius $a_{\mathrm{eff}}$ [^1] The effective grain radius gives the radius of the grain as if it were reshaped into a perfect sphere. from $50\,\mathrm{nm}$ up to $250\,\mathrm{nm}$ in steps of $50\,\mathrm{nm}$ were created. To reach an exact effective radius, the sizes of the final monomers added to the grain were sampled in a biased way (we lowered the upper bound on the monomer size), until the desired $a_{\mathrm{eff}}$ was reached exactly. In addition to BA, where each monomer remains at its initial sticking position, we modeled different monomer densities or porosities of aggregates by means of monomer migration. After hitting the forming grain, the added monomer rolls along the surface of the monomer it impacted until it touches a second monomer of the grain (BAM1). Here, all monomers are connected to at least two other monomers. For the BAM2 case each monomer rolls twice until it touches a third monomer. For BAM2, all monomers are connected to at least three other monomers. This is the densest way in which a random grain can be packed with monomers using the migration process. It follows that BA grains are the least dense/most porous and BAM2 grains are the most dense/least porous, an exemplary grain of each monomer density is shown in Fig. 1.
@@
To recapitulate, the three grain properties are the material from which the grain is made (astrosilicate or hydrogenated amorphous carbon), which differs in the optical properties; the monomer density of the grain (BA, BAM1, or BAM2), which determines how closely the monomers are packed; and finally, the grain size given by the effective radius ($50\,\mathrm{nm}$ to $250\,\mathrm{nm}$ in steps of $50\,\mathrm{nm}$). For every combination of these three properties, 30 different grains were created, leading to a total of 900 individual grains. This method of modeling ballistic grain aggregates was previously used in Reissl et al. (2023).
@@
Figure 1: Exemplary BA (left), BAM1 (middle), BAM2 (right) grains with an effective radius of ${a_{\mathrm{eff}}=250\ \mathrm{nm}}$. The rotation axis $\hat{a}_{\mathrm{1}}$ is depicted as the red arrow and is associated with the direction of the maximum moment of inertia $I_{\mathrm{max}}$.
3 Maximum grain rotation
@@
In the presence of a directed radiation field, irregularly shaped dust grains can experience momentum transfer from photon interactions. In general, the time-averaged forces acting on such a grain may result in a net torque, accelerating the rotation. This radiative torque (RAT) can be evaluated by (see e.g. Draine & Weingartner 1996, and LH07)
@@
where $\gamma$ describes the anisotropy of the radiation field, $\lambda$ is the wavelength, and the quantity $u_{\lambda}=F_{\lambda}/c$ is the spectral energy density of a radiation field with the spectral flux $F_{\lambda}$, whereas $c$ is the speed of light. These factors determine the overall strength of the torque, independent of the grain shape and orientation.
The effect of the individual grain shape and material is contained within the dimensionless spin-up efficiency vector ${Q}_{\Gamma}$ of the RAT. A perfectly spherical grain would experience no net torque at all because of the symmetry between collisions causing clockwise and anticlockwise torques. However, for grains with irregular shapes, the contributions of individual radiation-dust interactions lead to a nonzero ${Q}_{\Gamma}$ vector, which modulates the orientation, that is, the sign as well as the magnitude of ${\Gamma}_{\mathrm{RAT}}$.
In previous works, a power-law parameterization for $Q_{\Gamma}=|\@vec{Q}_{\Gamma}|$ was suggested (LH07)
@@
We refer to this relation as the canonical parameterization. Here, the constant part transitions into a power law at wavelengths equal to $1.8$ times the effective radius of the grain. The power-law exponent typically lies between $\alpha=-2.6$ (Herranen et al. 2019) and $\alpha=-4$ (LH07).
For our grain dynamics modeling, we assumed a stable alignment of the grain rotation with the direction of incoming radiation. For a single grain that stably rotates about its principal axis, the maximum possible angular momentum $\omega_{\mathrm{RAT}}$ caused by RATs can be calculated as
@@
where $\tau_{\mathrm{drag}}$ is the drag timescale, and $I_{\mathrm{max}}$ is the highest moment of inertia, defining the $a_{1}$-axis in the body-centric system of the grain (see e.g. Draine & Weingartner 1996, as well as Fig.  1). The total drag acting on the grain is composed of the collisions with the surrounding gas and the loss of angular momentum by emitting infrared photons. The total drag timescale can be evaluated from the gas-drag timescale $\tau_{\mathrm{gas}}$ and photon-drag timescale $\tau_{\mathrm{IR}}$ with
$\tau_{\mathrm{drag}}=\frac{1}{\tau_{\mathrm{gas}}^{-1}+\tau_{\mathrm{IR}}^{-1}},$ (4)
where a shorter timescale results in a stronger drag. Here, the gas-drag timescale
@@
The photon-drag timescale
$\tau_{\mathrm{IR}}=\frac{c^{2}I_{\mathrm{max}}}{a_{\mathrm{eff}}^{2}f_{\mathrm{IR}}(T_{\mathrm{dust}})}\quad$ (6)
depends on the integral over the photon emission,
@@
Here, $B_{\lambda}\left(T_{\mathrm{dust}}\right)$ is the Planck function modulated by the dimensionless efficiency of the absorption $Q_{\mathrm{abs}}(\lambda)$, and $T_{\mathrm{dust}}$ is the dust temperature. For simplicity, the dust is illuminated by a single source without extinction. For the dust heating, we assumed thermal equilibrium. Therefore, the temperature of a single dust grain can be evaluated via
$\int_{0}^{\infty}F_{\lambda}Q_{\mathrm{abs}}(\lambda)\,\mathrm{d}\lambda=4\pi\int_{0}^{\infty}B_{\lambda}\left(T_{\mathrm{dust}}\right)Q_{\mathrm{abs}}(\lambda)\,\mathrm{d}\lambda\,$ (8)
by equating the source flux $F_{\lambda}$ that is absorbed by the grain with the characteristic emitted flux $B_{\lambda}Q_{\mathrm{abs}}(\lambda)$ of each grain.
@@
 [^2] The table includes the stellar radius $R$ and the effective temperature $T_{\mathrm{eff}}$ for stellar sources S1 and S2, the bolometric luminosity $L_{\mathrm{bol}}$ for all localized sources, the anisotropy $\gamma$, and the radiation energy density $u$, given in relation to $u_{\mathrm{ISRF}}=8.64\cdot 10^{-14}\ \mathrm{J\ m^{-3}}$. The distance to the localized sources S1, S2, and to the AGN is taken to be $10\,\mathrm{pc}$.
Figure 2: Spectral flux of sources ISRF, S1, S2, and AGN that were used to irradiate the grain aggregates.